# Benchmarking Linguistic Diversity of Large Language Models

### v1

@@ Abstract
The development and evaluation of Large Language Models (LLMs) has primarily focused on their task-solving capabilities, with recent models even surpassing human performance in some areas. However, this focus often neglects whether machine-generated language matches the human level of diversity, in terms of vocabulary choice, syntactic construction, and expression of meaning, raising questions about whether the fundamentals of language generation have been fully addressed. This paper emphasizes the importance of examining the preservation of human linguistic richness by language models, given the concerning surge in online content produced or aided by LLMs. We propose a comprehensive framework for evaluating LLMs from various linguistic diversity perspectives including lexical, syntactic, and semantic dimensions. Using this framework, we benchmark several state-of-the-art LLMs across all diversity dimensions, and conduct an in-depth case study for syntactic diversity. Finally, we analyze how different development and deployment choices impact the linguistic diversity of LLM outputs.
@@ 1 Introduction
Recent studies have highlighted concerns regarding the linguistic diversity of LLM outputs. By comparing human and model-generated content, researchers have shown that models frequently struggle to reflect the nuances and variations characteristic of human expression (Shaib et al., 2024a; Giulianelli et al., 2023). Additionally, these concerns are reinforced by findings that training language models on synthetic text can lead to a further decline in linguistic diversity (Guo et al., 2024b).
In fact, LLMs tend to be inherently conservative in producing diverse content. During training, models undergo homogenization to the most frequent patterns in the training data, where creative outlier narratives, views, styles, and knowledge are often underrepresented (Kandpal et al., 2023). Unlike models, human language production involves a complex interplay of factors that go beyond merely optimizing probabilities (Holtzman et al., 2020). It is therefore crucial to emphasize evaluating output diversity in language models and systematically consider these metrics to guide future model development and deployment decisions.
Currently, a principled and comprehensive evaluation framework for linguistic diversity is lacking in the literature (Shaib et al., 2024a). While some studies on Natural Language Generation (NLG) report diversity metrics, they typically focus on a single diversity aspect (e.g., lexical diversity (Chakrabarty et al., 2022)), often experimenting within a single domain and task (e.g., news summarization  (Shaib et al., 2024a)). This narrow focus is problematic since diversity varies across aspects and depends on the domain (Guo et al., 2024b). Although some efforts have been made to assess the influence of reinforcement learning from human feedback (RLHF) on diversity (Kirk et al., 2024), the impact of other key design and development stages—such as model scale, quantization, and decoding strategy—remains unexplored. Additionally, there is a limited understanding of how LLMs develop the capability to generate diverse language through successive pretraining checkpoints. Ultimately, no study has benchmarked the diversity performance of state-of-the-art LLMs across different aspects and domains.
In this work, we first establish a framework for evaluating linguistic diversity of LLM outputs on a corpus level. We then benchmark six prominent LLMs on five NLG tasks, and compare the diversity of their outputs across three different aspects: lexical, syntactic, and semantic. We further explore syntactic diversity, and conduct a case study comparing the distribution of dependency trees generated by LLMs and humans. Finally, we investigate how LLM output diversity changes across different development stages, and with varying decisions of deployment. The main research questions we address are as follows:
@@
    How do different design (e.g., model scale, training data) and deployment (e.g., decoding strategy, quantization) choices affect diversity? (See § 6.2 and § 6.3)
It is worth noting that we study linguistic diversity in a monolingual context, focusing on the English language. However, the evaluation methodology is language agnostic and could easily be extended to other languages, given that employed NLP toolkits (e.g., dependency parsers, sentence embeddings) exist for the language. Furthermore, our approach to analyzing the influence of various factors on LLM outputs is adaptable to other dimensions, such as linguistic naturalness (Guo et al., 2024a). The code is publicly available [^1] https://github.com/YanzhuGuo/llm-diversity.
@@ 2.1 Evaluation of Human Language
The evaluation of linguistic diversity in model-generated language has emerged as a relatively recent focus of research. This development is driven, in part, by growing concerns over the increasing prevalence of model-generated or model-influenced content online, prompting questions about whether LLMs can maintain the linguistic richness characteristic of human language (Guo et al., 2024b). Moreover, advances in language generation quality have brought model outputs closer than ever to human-level coherence and plausibility. Assessing linguistic diversity becomes meaningful only when the generated text meets these standards. For example, a randomly initialized model might produce token sequences with high lexical diversity, but such outputs hold no practical value (Uchendu et al., 2023).
@@ 2.2 Evaluation of Generated Language
To the best of our knowledge, Tevet and Berant (2021) were the first authors to systematically evaluate diversity in NLG. They proposed to create diversity metrics from any two-sentence similarity measure, defining diversity as the inverse of the mean similarity score across all unordered pairs. N-gram-based metrics were used to assess form diversity, while model-based metrics like Sentence-BERT similarity measured content diversity. They concluded that a notable disparity exists between automatic metrics and human judgment, and that human evaluation of diversity becomes challenging in sets with more than ten responses. Since then, additional metrics have been proposed to capture linguistic diversity, including semantic diversity metrics based on natural language inference (Stasaski and Hearst, 2022) or semantic entropy (Han et al., 2022), and syntactic diversity metrics derived from n-grams of Part-of-Speech (POS) tags (Giulianelli et al., 2023) or graph similarity kernels of syntax trees (Guo et al., 2024b).
@@ 2.3 Impact of LLMs on Linguistic Diversity
Diverging from the above research focused on developing methods to evaluate linguistic diversity, another line of work explores the the impact of LLMs on both human and model generated text, often demonstrating a decline in diversity. Guo et al. (2024b) showed that iteratively training LLMs on synthetic data generated by earlier models, leads to a consistent decline in lexical, syntactic, and semantic diversity, especially for tasks requiring high creativity. Similarly, Padmakumar and He (2024) reported a statistically significant reduction in linguistic diversity when humans write with InstructGPT. This reduction in linguistic diversity is also observed in other contexts: Liang et al. (2024) identified a significant frequency shift toward LLM-preferred words in academic writing, and Luo et al. (2024) reported reduced morphosyntactic diversity in machine translations compared to human translations.
Closely related to our work, Kirk et al. (2024) examined how SFT and preference tuning affect LLM generalization and diversity. They found that preference tuning substantially reduces lexical and semantic diversity compared to SFT. Our research also explores the factors that influence diversity while broadening the analysis to include a wider range of diversity aspects, models, tasks and factors. Moreover, our findings on the impact of preference tuning differ from those of Kirk et al. (2024), likely due to differences in task domain, accentuating the importance of contextualizing conclusions.
@@ 3 Metrics for Linguistic Diversity
According to Tevet and Berant (2021), diversity can be viewed as having a hierarchical structure with various sub-aspects. Lexical diversity and syntactic diversity are considered sub-aspects of form diversity, while semantic diversity reflects content diversity. Although there are potentially other sub-aspects of linguistic diversity, such as style (register) diversity within form diversity, these elements are often more ambiguous, harder to quantify, and tend to overlap with the existing sub-aspects. For example, style diversity is partially captured through lexical and syntactic diversity, as they reflect preferences in vocabulary and grammatical structures. Consequently, we focus on the three aspects of diversity that are more clearly defined, easier to quantify, and have relatively low correlation with each other (further discussed in Section 5.1).
@@
In our study, we choose to measure across-input diversity, as we focus on linguistic patterns across a broad range of generations. Formally, given a set of generated outputs $S=\{s_{1},s_{2},\dots,s_{n}\}$, we compute $Div(S)$ differently depending on the aspect of diversity: for lexical diversity, $S$ is treated as a set of n-grams, while for syntactic and semantic diversity, $S$ is considered as a set of sentences.
In the following sections, we explain each diversity aspect and the specific metrics used.
@@ 3.1 Lexical Diversity
We employ Unique-$n$ (Johnson, 1944; Templin, 1957), established for evaluating lexical diversity. It is calculated as the ratio of unique $n$-grams to the total number of $n$-grams. When $n=1$, it is equivalent to Type-Token Ratio (Johnson, 1944; Templin, 1957). We report the average Unique-$n$ across unigrams, bigrams, and trigrams. Originally used in child language research, Unique-$n$ is useful for assessing language development, where a lower value might indicate limited lexical variety (Miller, 1981). We use the global Unique-$n$ measure rather than the moving average Unique-$n$ because we are interested in the overall diversity capabilities of LLMs across different inputs rather than their performance on individual inputs. Moving average methods might miss global lexical repetitions due to their localized nature (Bestgen, 2023). To mitigate the influence of output length on Unique-$n$, we always randomly choose 40K samples to constitute the set of $n$-grams for each $n$.
Figure 1: Lingustic diversity benchmarking results for NLG tasks detailed in Table 1.
Dataset Input Output
Language Modeling (LM) Wikitext-2 (Merity et al., 2017) Block of 128 tokens from Wikipedia Prediction of the next block
Machine Translation (MT) WMT-14 (Bojar et al., 2014) News story sentence in French Corresponding English translation
Summarization (Summ) XLSUM (Hasan et al., 2021) Full news article from the BBC Summary of the news article
Next Utterance Generation (NUG) DailyDialog (Li et al., 2017) Scripted dialogue on general topics Next utterance of the dialogue
Automatic Story Generation (ASG) WritingPrompts (Fan et al., 2018) Story prompt shared by Reddit users Story continuation based on the prompt
Table 1: Summary of datasets, inputs, and outputs for benchmarked NLG tasks.
@@ 3.1.1 Syntactic Diversity
To address this limitation, we employ a graph-based metric for quantifying syntactic diversity (Guo et al., 2024b). This metric relies on a neural parser (Qi et al., 2020) to generate dependency trees from sentences, following the universal dependencies framework. These trees are converted into graph representations, where nodes represent words and edges denote dependency relationships. The nodes are labeled by the PoS tag of each word. The Weisfeiler-Lehman (WL) graph kernel (Shervashidze et al., 2011; Siglidis et al., 2020) is then applied to map these graphs into a vector space. This kernel, based on the WL isomorphism test, positions structurally similar graphs closer together in the vector space. Syntactic diversity is then measured using the average pairwise distance between graphs, formalized as: $\text{Div}_{\text{syn}}(S)=\frac{1}{\binom{n}{2}}\sum_{1\leq i<j\leq n}WL(s_{i},s_{j})$. Alternatively, the pairwise distances between the dependency trees could also be measured by the tree editing distance (Zhang and Shasha, 1989). However, the algorithm to compute the tree editing distance has much higher computational complexity and is thus not scalable to a large set of texts.
@@ 3.1.2 Semantic Diversity
Semantic diversity refers to the range and variety of meanings or ideas conveyed within a text or set of texts. It evaluates how broadly and uniquely different concepts, topics, or ideas are expressed, reflecting the depth and scope of the content. High semantic diversity suggests a text covers diverse ideas or meanings, while low semantic diversity implies repetition or a narrow focus on specific concepts. Recent studies (Tevet and Berant, 2021; Stasaski and Hearst, 2022) have pointed out that traditional lexical metrics may not fully capture semantic diversity. Similar words can convey different meanings, and different words can convey similar meanings (Yarats and Lewis, 2018).
To address this, we first convert sentences into semantically meaningful embeddings using Sentence-BERT Reimers and Gurevych (2019). Semantic diversity is then quantified as the dispersion of these embeddings in the semantic space, measured by the average pairwise cosine distance (scaled to the range $[0,1]$) between all embedding vectors: $\text{Div}_{\text{sem}}(S)=\frac{1}{\binom{n}{2}}\sum_{1\leq i<j\leq n}\frac{1+d_{\text{cos}}}{2}(e(s_{i}),e(s_{j}))$, where $e$ represents Sentence-BERT embeddings.
@@ 4 Settings for Diversity Benchmarking
We outline the tasks, datasets, and models used to establish our linguistic diversity benchmark.
Generation tasks. To effectively compare the linguistic diversity of LLM outputs across various scenarios, we selected five tasks with progressively increasing levels of “creativity”: language modeling, machine translation, summarization, automatic story generation, and next utterance generation. Table 1 outlines the inputs, outputs, and datasets associated with each task. Note that the input is combined with a task-specific instruction (e.g., “Continue the following story:” for story generation) to construct the final prompt.
For each task, we randomly select 10K samples from the original dataset. To obtain the benchmark results, we decode the outputs for all models and tasks using a combination of nucleus sampling (t=0.6) and top-k sampling (k=0.9). We further analyze in Section 6.2 the impact of different decoding parameters on output diversity.

### v2

@@ Abstract
The development and evaluation of Large Language Models (LLMs) has primarily focused on their task-solving capabilities, with recent models even surpassing human performance in some areas. However, this focus often neglects whether machine-generated language matches the human level of diversity, in terms of vocabulary choice, syntactic construction, and expression of meaning, raising questions about whether the fundamentals of language generation have been fully addressed. This paper emphasizes the importance of examining the preservation of human linguistic richness by language models, given the concerning surge in online content produced or aided by LLMs. We adapt a comprehensive framework for evaluating LLMs from various linguistic diversity perspectives including lexical, syntactic, and semantic dimensions. Using this framework, we benchmark several state-of-the-art LLMs across all diversity dimensions, and conduct an in-depth analysis for syntactic diversity. Finally, we analyze how the design, development and deployment choices of LLMs impact the linguistic diversity of their outputs, focusing on the creative task of story generation.
@@ 1 Introduction
Recent studies have highlighted concerns regarding the linguistic diversity of LLM outputs. By comparing human and model-generated content, researchers have shown that models frequently struggle to reflect the nuances and variations found in human expression (Shaib et al., 2024a; Giulianelli et al., 2023). Additionally, these concerns are reinforced by findings that training language models on synthetic text can lead to a further decline in linguistic diversity (Guo et al., 2024b).
In fact, LLMs tend to be inherently conservative in producing diverse content. During training, models undergo homogenization to the most frequent patterns in the training data, where creative outlier narratives, views, styles, and knowledge are often underrepresented (Kandpal et al., 2023). Unlike models, human language production involves a complex interplay of factors that go beyond merely optimizing probabilities (Holtzman et al., 2020). It is therefore crucial to emphasize evaluating output diversity in language models and systematically consider these metrics to guide future model design, development and deployment decisions.
Currently, principled and comprehensive studies on evaluating linguistic diversity are lacking in the literature (Shaib et al., 2024a). While some works on Natural Language Generation (NLG) report diversity metrics, they typically focus on a single diversity aspect (e.g., lexical diversity (Chakrabarty et al., 2022)), often experimenting within a single domain and task (e.g., news summarization  (Shaib et al., 2024a)). This narrow focus is problematic since diversity varies across aspects and depends on the domain (Guo et al., 2024b). Although some efforts have been made to assess the influence of reinforcement learning from human feedback (RLHF) on diversity (Kirk et al., 2024), the impact of other key design and development stages—such as model scale, quantization, decoding strategy, and prompt formulation—remains unexplored. Additionally, there is a limited understanding of how LLMs develop the capability to generate diverse language through successive pretraining checkpoints. Ultimately, no study has benchmarked the diversity performance of state-of-the-art LLMs across different aspects and domains.
In this work, we first establish a framework for evaluating linguistic diversity of LLM outputs on a corpus level. We then benchmark six prominent LLMs on five NLG tasks, and compare the diversity of their outputs across three different aspects: lexical, syntactic, and semantic. We place particular emphasis on story generation, the most creative task where linguistic diversity plays a crucial role, conducting a deeper analysis in this context. Specifically, we examine syntactic diversity through a case study comparing the distribution of dependency trees in human-written and LLM-generated texts. Finally, we also investigate how LLM output diversity changes across different development stages, and with varying decisions of deployment.
The main research questions we address are as follows:
@@
    How do different design (e.g., model scale, training data) and deployment (e.g., decoding strategy, prompt formulation, quantization) choices affect diversity? (See § 6.2, § 6.3 and § 6.4)
It is worth noting that we study linguistic diversity in a monolingual context, focusing on the English language. However, the evaluation methodology is language agnostic and could easily be extended to other languages, given that employed NLP toolkits (e.g., dependency parsers, sentence embeddings) exist for the language. Furthermore, our approach to analyzing the influence of various factors on LLM outputs is adaptable to other dimensions, such as linguistic naturalness (Guo et al., 2024a).
The code for our research is available at https://github.com/YanzhuGuo/llm-diversity.
@@ 2.1 Evaluation of Human Language
The evaluation of linguistic diversity in model-generated language has emerged as a relatively recent focus of research. This development is driven, in part, by growing concerns over the increasing online prevalence of model-generated or model-influenced content (Geng and Trotta, 2024), prompting questions about whether LLMs can reflect the linguistic richness characteristic of human language (Guo et al., 2024b). However, assessing linguistic diversity is meaningful only when the generated text meets basic standards of quality. For instance, a randomly initialized model might produce token sequences with high lexical diversity, but such outputs lack any practical value (Uchendu et al., 2023). Recent advances in language generation quality have brought model outputs closer than ever to human-like coherence and plausibility, making the evaluation of linguistic diversity more relevant and necessary than before.
@@ 2.2 Evaluation of Generated Language
To the best of our knowledge, Tevet and Berant (2021) were the first authors to systematically evaluate diversity in NLG. They proposed to create diversity metrics from any two-sentence similarity measure, defining diversity as the inverse of the mean similarity score across all unordered pairs. N-gram-based metrics were used to assess form diversity, while model-based metrics like Sentence-BERT similarity measured content diversity. They concluded that a notable disparity exists between automatic metrics and human judgment, and that human evaluation of diversity becomes challenging in sets with more than ten sentences.
Since then, additional metrics have been proposed to capture linguistic diversity, including semantic diversity metrics based on natural language inference (Stasaski and Hearst, 2022) or semantic entropy (Han et al., 2022), and syntactic diversity metrics derived from n-grams of Part-of-Speech (POS) tags (Giulianelli et al., 2023) or graph similarity kernels of syntax trees (Guo et al., 2024b).
Another relevant research direction involves divergence-based metrics that compare the distributions of human-written and machine-generated text. Examples included MAUVE (Pillutla et al., 2021), which leveraged distributions of GPT-2 embeddings, as well as later approaches based on specific linguistic features (Guo et al., 2024a). While such metrics do not explicitly measure linguistic diversity, they can offer insights into distributional differences, of which diversity is a key component.
@@ 2.3 Impact of LLMs on Linguistic Diversity
Diverging from the above research focused on developing methods to evaluate linguistic diversity, another line of work explores how LLM-generated content impact future models or human writing patterns, often demonstrating a decline in diversity. Guo et al. (2024b) showed that iteratively training LLMs on synthetic data generated by earlier models leads to a consistent decline in lexical, syntactic, and semantic diversity, especially for tasks requiring high creativity. Similarly, Padmakumar and He (2024) reported a statistically significant reduction in linguistic diversity when humans write with InstructGPT. This reduction in linguistic diversity is also observed in other contexts: Liang et al. (2024) identified a significant frequency shift toward LLM-preferred words in academic writing, while Luo et al. (2024) reported reduced morphosyntactic diversity in machine translations compared to human translations.
Closely related to our work, Kirk et al. (2024) examined how SFT and preference tuning affect LLM generalization and diversity. They found that preference tuning substantially reduces lexical and semantic diversity compared to SFT. Our research also explores the factors that influence diversity while broadening the analysis to include a wider range of diversity aspects, models, tasks and factors. However, our findings on the impact of preference tuning differ from those of Kirk et al. (2024), likely due to differences in task domain, accentuating the importance of contextualizing conclusions.
@@ 3 Metrics for Linguistic Diversity
According to Tevet and Berant (2021), diversity can be divided into two primary dimensions: form diversity and content diversity. Lexical and syntactic diversity are sub-aspects within form diversity, whereas semantic diversity pertains to content diversity. While additional sub-aspects of form diversity, such as style diversity, exist and could potentially be measured through style representations (Soto et al., 2024), these aspects are generally less interpretable and often overlap with other dimensions of diversity. For instance, style diversity inherently intersects with lexical and syntactic diversity, as stylistic choices typically involve preferences in vocabulary and grammar. Therefore, in this study, we concentrate on the three diversity aspects (lexical, syntactic, and semantic) that are clearly defined, straightforward to interpret, and exhibit relatively low mutual correlation (further detailed in Section 5.1).
@@
In our study, we choose to measure across-input diversity, as we focus on general linguistic patterns across a broad range of generations. Formally, given a set of generated outputs $S=\{s_{1},s_{2},\dots,s_{n}\}$, we compute $Div(S)$ differently depending on the aspect of diversity: for lexical diversity, $S$ is treated as a set of n-grams, while for syntactic and semantic diversity, $S$ is considered as a set of sentences.
We build on the linguistic diversity evaluation framework and preprocessing methods of Guo et al. (2024b), but shift the focus from studying the effects of recursive synthetic training on OPT (Zhang et al., 2022) to comparing linguistic diversity across a range of state-of-the-art LLMs. We also investigate how various design choices such as model scale and training data, and deployment factors such as decoding strategy, prompt formulation and quantization, impact diversity. In principle, the same research protocol could be extended to examine per-input diversity, allowing for the investigation of uncertainty and variability in text generation (Giulianelli et al., 2023). Although this lies beyond the scope of the current study, it represents a promising direction for future work.
In the following sections, we describe each aspect of diversity and the specific metrics used to assess them.
Instruction Input Output
Language Modeling (LM) Not applicable (no instruction) Block of 128 tokens from Wikipedia Prediction of the next block
Machine Translation (MT) Translate from French to English News story sentence in French Corresponding English translation
Summarization (Summ) Summarize the following article Full news article from the BBC Summary of the news article
Next Utterance Generation (NUG) Continue the following dialogue Scripted dialogue on general topics Next utterance of the dialogue
Automatic Story Generation (ASG) Continue the following story Story prompt shared by Reddit users Story continuation based on the prompt
Table 1: Summary of instructions, inputs, and outputs for benchmarked NLG tasks. Task instructions are placed in the system input when supported, otherwise prepended to the user input.
@@ 3.1 Lexical Diversity
[…]
@@ 3.1.1 Syntactic Diversity
To address this limitation, we employ a graph-based metric for quantifying syntactic diversity (Guo et al., 2024b). This metric relies on a neural parser (Qi et al., 2020) to generate dependency trees from sentences, following the universal dependencies framework. In these trees, nodes represent words and edges capture syntactic dependencies, with nodes labeled by the corresponding part-of-speech (PoS) tags. The Weisfeiler-Lehman (WL) graph kernel (Shervashidze et al., 2011; Siglidis et al., 2020) is applied to map these trees into a reproducing kernel Hilbert space, where structurally similar graphs are positioned closer together based on the WL isomorphism test. Syntactic diversity is then computed as the average pairwise distance between these graphs, formalized as: $\text{Div}_{\text{syn}}(S)=\frac{1}{\binom{n}{2}}\sum_{1\leq i<j\leq n}WL(s_{i},s_{j})$.
@@ 3.1.2 Semantic Diversity
Semantic diversity refers to the range and variety of meanings or ideas conveyed within a text or set of texts. It evaluates how broadly and uniquely different concepts, topics, or ideas are expressed, reflecting the depth and scope of the content. Low semantic diversity often indicates repetition or a narrow focus, whereas high semantic diversity typically suggests coverage of a wide array of topics. However, texts should meet a basic quality standard before being evaluated for semantic diversity, since high semantic diversity can also arise from noisy or irrelevant content. Recent studies (Tevet and Berant, 2021; Stasaski and Hearst, 2022) have pointed out that traditional lexical metrics may not fully capture semantic diversity. Similar words can convey different meanings, and different words can convey similar meanings (Yarats and Lewis, 2018).
To address this, we first convert sentences into semantically meaningful embeddings using Sentence-BERT Reimers and Gurevych (2019). Semantic diversity is then quantified as the dispersion of these embeddings in the semantic space, measured by the average pairwise cosine distance (scaled to the range [0,1][0,1][ 0 , 1 ]) between all embedding vectors: $\text{Div}_{\text{sem}}(S)=\frac{1}{\binom{n}{2}}\sum_{1\leq i<j\leq n}\frac{d_{\text{cos}}(e(s_{i}),e(s_{j}))}{2}$, where $e$ represents Sentence-BERT embeddings.
Figure 1: Linguistic diversity benchmarking results for NLG tasks detailed in Table 1.
@@ 4 Settings for Diversity Benchmarking
We outline the tasks and models used to establish our linguistic diversity benchmark. We decode the outputs for all tasks and models using a combination of nucleus sampling (t=0.6) and top-k sampling (k=0.9). We further analyze in Section 6.2 the impact of different decoding parameters on output diversity.
4.1 Generation Tasks
To effectively compare the linguistic diversity of LLM outputs across various scenarios, we choose five tasks with progressively increasing levels of “creativity”. The inputs, outputs, and instructions for each task are summarized in Table 1. To maintain general model behavior and avoid overly influencing responses through prompt design, we keep the instructions minimal.
For each task, we randomly select 10K samples from the original dataset. While our experiments are conducted using a single dataset per task, we deliberately select the most representative for each. Nonetheless, the conclusions drawn from our experiments should be interpreted within the context of these datasets. We now provide a detailed introduction to each task and its associated dataset.
Language Modeling (LM) involves predicting the next token in a sequence based on the preceding tokens and is fundamental to all NLG applications. We use the Wikitext-2 dataset (Merity et al., 2017) to evaluate general purpose language modeling. Derived from Wikipedia articles, Wikitext-2 offers a rich corpus with around 2 million tokens across diverse topics. We chunk texts into blocks of 128 tokens and ask the models to predict the next 128 tokens. Language modeling serves as the basis of all other tasks, so it is considered as the least creative.
Machine translation (MT) aims to transfer text from one language to another while maintaining the original meaning. We use the WMT-14 dataset (Bojar et al., 2014) which contains parallel corpora for multiple language pairs. For our experiments, we focus on a subset of this benchmark that includes French-to-English sentence pairs from multiple sources. We classify this task as having a low level of creativity, as the output is expected to convey exactly the same meaning as the input.
Summarization (Summ) is the process of generating concise summaries of lengthy texts, preserving key information while minimizing redundancy. We use the XLSUM dataset (Hasan et al., 2021), which features news articles in various languages along with their summaries. Our experiments focus on the English portion of the dataset. While we also categorize this task as low in creativity, it allows slightly more flexibility than machine translation, as the model must decide which information to prioritize and include in the summary.
Next Utterance Generation (NUG) aims to produce natural utterances in conversations while maintaining contextual relevance. For this task, we use the DailyDialog dataset (Sai et al., 2020), a human-curated multi-turn dialogue corpus designed to cover a broad range of topics relevant to everyday interactions. In our setup, the model is always prompted to predict the final utterance based on all preceding dialogue turns. We consider next utterance generation to be a creative task, as there is a large space of possible and coherent utterances in response to a certain dialog context. However, the everyday nature and structure of the dataset place some limits on the level of creativity.
Automatic Story Generation (ASG) centers on producing engaging and coherent narratives from story prompts or initial contexts. We employ the WritingPrompts dataset (Fan et al., 2018), which comprises prompts and corresponding stories contributed by Reddit users. It includes a wide variety of prompts in different formats, encouraging diverse and creative responses. Among our tasks, we consider story generation to be the most creative, as the prompts typically impose minimal constraints on narrative structure and content, allowing for maximal expressive freedom.